1. Your firm purchased equipment costing $600,000 1 year ago. The firm also incurred shipping costs of $18,000 and installation costs of $2,000. The IRS allows this category of equipment to be depreciated over 4 years. The equipment is expected to sell for $40,000 today. What is the book value of the asset at the time of the sale?

2. A piece of equipment was originally purchased 5 years ago for $400,000 and is now being sold for $320,000. Yearly depreciation of $30,000 has been recorded. Based on a gain of $70,000, $17,500 is now owed in taxes. What will be recorded as the net proceeds from this sale?

3. An asset was purchased for $300,000 originally. Each year for the past 8 year, depreciation has been recorded which results in the asset having a current book value of $180,000. Now your firm is selling the asset for $150,000. What will be the total tax burden (or tax credit) caused by this sale? If it is a tax burden, enter the number as a positive number. If it is a tax credit, enter the number as a negative number. Assume a tax rate of 40%.
Answer #1

1.
=(600000+18000+2000)*3/4=465000

2.
=320000-17500=302500

3.
We see that as sale price is less than book value, there will be tax credit=(150000-180000)*40%
=-12000.00